OpenStudy (anonymous):

Given a + b + c = 5 and a² + b² + c² = 9, find the value of ab + ac + bc

OpenStudy (anonymous):

ab+ac+bc=8

OpenStudy (anonymous):

procedure: \[(a+b+c)^2=a^2+b^2+c^2+2(ab+ac+bc)\] \[5^2=9+2(ab+ac+bc)\] \[ab+ac+bc=8\]